The main road in Belleville is 7 3/10 miles long. So far, 2 3/4 miles have been repaved. How many miles have not been repaved?
Answer by checkley71(8403)

7 3/10-2 3/4 OR 7.3-2.75=4.55 MILES TO BE PAVED. DECIMAL SOLUTION
FRACTION SOLUTION 7 3/10=73/10 & 2 3/4=11/4 SO 73/10-11/4 REQUIRES FINDING A COMMON DENOMINATOR THAT 4 & 10 is evenly divisable by. finding the factors or 4 & 10 helps find the lowest commmon denominator thus 2*2=4 & 2*5=10 both denominators have a common factor of 2 therefotre we only need one of the 2s but we need the other 2 & the 5 thus 2*2*5=20 which is the lowest common denominator. you could have just multiplied 4*10=40 but this would not be the lowest common denominator but a common denominator that works.
now change each fraction into a fraction with a denominator of 20 by multiplying the 73/10 by 2/2 we get 146/20 & multiplying 11/4 by 5/5 we get 55/20. note that 2/2 & 5/5 is just another way of expressing the number 1 thus any number multiplied by 1 is the same number just in a different form.
now we can subtract 55/20 from 146/20 thus (146-55)/20 or 91/20 or 4 11/20 & 11/20=.55 so we still have an answer of 4.55